# Continuous sentiment scores for literary and multilingual contexts

### v1

@@ 2.4 Models
The implementation of Concept Vector Projection used to classify sentiment in this paper is based on the language model paraphrase-multilingual-mpnet-base-V2121212https://huggingface.co/sentence-transformers/paraphrase-multilingual-mpnet-base-v2 [28]. This is a 278M parameter model, based on a mean-pooled BERT architecture, optimized for sentence similarity by using Siamese and Triplet networks. This model was chosen because of its multilingual capabilities and excellent size-to-performance ratio. Investigations during model selection indicate that a larger model may increase model correlation with human ratings in exchange for compute budget.
@@ 3.1 Continuous scoring
Figure 3: Scatterplot of Sentiment Predictions for respectively Sentiment Projection and xlm-roberta. While the xlm-roberta model, in theory, can predict a continuous space of sentiments when transforming it with confidence scores, visual inspection shows that certain ranges of the sentiments spectrum are not used. While both models achieve high correlations, it appears that xlm-roberta achieves this by matching human tendencies to predict neutral.
A key benefit of the Sentiment Projection model is its ability – like dictionary tools – to produce genuinely continuous predictions. In contrast, Transformer-based token-classification models such as xlm-roberta, which can be coerced to output continuous scores (see subsubsection 2.2.2), in practice exhibit a “pseudo-trinary” behavior: their predictions cluster heavily at zero and at the two polar extremes. This behavior is clearly visible both in the scatterplots of predicted vs true sentiments (Figure 3) and in the histograms of model outputs (Appendix A, Figure 5). When looking at the EmoBank results (Appendix A, Figure 4), the discretized output of xlm-roberta appears even more sharply tri-modal than the human scores, which average ten annotators.
@@ 3.2 Performance on literary data across genres
We evaluated all models on the full multilingual Fiction4 corpus. For the dictionary-based tools (VADER and Syuzhet), originally Danish texts were translated into English (see subsubsection 2.2.1). Danish-specific models generally underperform on genres that are natively English (Prose, Poetry), which drags down their overall correlation scores. An outlier is danish-sentiment, which delivers relatively consistent results across both languages; however, it still falls short of MeMo-BERT-SA on the original Danish texts – most notably in the Fairy Tales genre.
@@ 3.4 Performance on literary and non-literary contemporary data
To make sure that our model does not overfit its sentiment vector to the in-context sentiment cues of the stories in the Fiction4 corpus, we tested it against the EmoBank dataset – which consists of contemporary literary and non-literary data. All Multilingual and dictionary-based models were also tested for their correlation with the human gold standard of the EmoBank dataset. The Sentiment Projection Model still achieved the highest overall correlation with human ratings. Although it shows a lower correlation for a few genres (i.a., Letters), it still appears to generalize well to contemporary out-of-training distribution data. It should be noted that the model outperforms the other models the most in the fiction genre, indicating that the sentiment vector may be slightly fine-tuned or overfit to fiction-specific sentiment indicators. [^13] While this can also be a drawback, it supports the idea that domain-specific sentiment analysis can be highly beneficial. For example, a sentiment analysis method for fiction should be sensitive to the specific omissive, implicit, and concrete object-based sentiment cues that are rarer in other genres [3]. [14] suggests that travel guides use similar mechanisms – sentiment is evoked through unsentimental, descriptive, and concrete detail. The fact that Sentiment Projection performs well also for both genres suggests it captures this kind of indirect sentiment expression.
@@ 4 Discussion & conclusions
While this paper corroborates the findings of [13], showing that translation (even without a quality check) to English increases the similarity of human and transformer-model scores, it also shows that this is not the case for Sentiment Projection, which performs slightly better on the original (Danish) sentences.
Finally, the workflow presented in Figure 2 has been used to design a sentiment model, but allows easy generalization to other concepts of choice. The method could also work for other emotional concepts, such as emotion recognition, language detection, or abstract concepts like a nature-to-industry gradient. We encourage curious readers to search for inspiration for potential vectors in Linear Representation Hypothesis [25] and Steering Vector [35] literature. Due to the flexible nature of the algorithm, there is no rigid lower boundary on the number of training points required for a stable vector, although the chances of over-representing the non-concept context of training sentences naturally increase as the number of sentences decreases. A future empirical investigation of the stability of the vector when using smaller training sets would be a useful guide.

### v2

@@ 2.4 Models
The implementation of Concept Vector Projection used to classify sentiment in this paper is based on the language model paraphrase-multilingual-mpnet-base-V2121212https://huggingface.co/sentence-transformers/paraphrase-multilingual-mpnet-base-v2 [reimers-2019-sentence-bert]. This is a 278M parameter model, based on a mean-pooled BERT architecture, optimized for sentence similarity by using Siamese and Triplet networks. This model was chosen because of its multilingual capabilities and excellent size-to-performance ratio. Investigations during model selection indicate that a larger model may increase model correlation with human ratings in exchange for compute budget.
@@ 3.1 Continuous scoring
Figure 3: Scatterplot of Sentiment Predictions for respectively Sentiment Projection and xlm-roberta. While the xlm-roberta model, in theory, can predict a continuous space of sentiments when transforming it with confidence scores, inspection shows that certain ranges of the sentiments spectrum are not used. While both models achieve high correlations, it appears that xlm-roberta achieves this by matching human tendencies to predict neutral.
A key benefit of the Sentiment Projection model is its ability – like dictionary tools – to produce genuinely continuous predictions. In contrast, Transformer-based token-classification models such as xlm-roberta, which can be coerced to output continuous scores (see subsubsection 2.2.2), in practice exhibit a “pseudo-trinary” behavior: their predictions cluster heavily at zero and at the two polar extremes. This behavior is visible both in the scatterplots of predicted vs true sentiments (Figure 3) and in the histograms of model outputs (Appendix A, Figure 5). When looking at the EmoBank results (Appendix A, Figure 4), the discretized output of xlm-roberta appears even more sharply tri-modal than the human scores, which average ten annotators.
@@ 3.2 Performance on literary data across genres
We evaluated all models on the full multilingual Fiction4 corpus. For the dictionary-based tools (VADER and Syuzhet), originally Danish texts were translated into English (see subsubsection 2.2.1). Danish-specific models generally under-perform on genres that are (originally) in English (Prose, Poetry), which drags down their overall correlation scores. An outlier is danish-sentiment, which delivers relatively consistent results across both languages; however, it still falls short of MeMo-BERT-SA on the original Danish texts – most notably in the Fairy Tales genre.
@@ 3.4 Performance on literary and non-literary contemporary data
To make sure that our model does not overfit its sentiment vector to the in-context sentiment cues of the stories in the Fiction4 corpus, we tested it against the EmoBank dataset – which indexes contemporary literary and non-literary data. All Multilingual and dictionary-based models were tested for their correlation with the human gold standard of the EmoBank dataset. The Sentiment Projection Model still achieved the highest overall correlation with human ratings. Although it shows a lower correlation for a few genres (i.a., Letters), it still appears to generalize well to contemporary out-of-training distribution data. It should be noted that the model outperforms the other models the most in the fiction genre, indicating that the sentiment vector may be slightly fine-tuned or overfit to fiction-specific sentiment indicators. While this can also be a drawback, it supports the idea that domain-specific sentiment analysis can be highly beneficial. For example, a sentiment analysis method for fiction should be sensitive to the specific sentiment cues (like omission, implicitness, concrete and object-based, etc.), rarer in other genres [bizzoni_feldkamp_sea_2024]. [feldkamp2024sentiment] suggests that travel guides use similar mechanisms – sentiment is evoked through unsentimental, descriptive, and concrete detail. The fact that Sentiment Projection performs well also for both genres suggests it captures this kind of indirect sentiment expression.
@@ 4 Discussion & conclusions
While this paper corroborates the findings of [feldkamp_comparing_2024], showing that translation (even without a quality check) to English increases the similarity of human and transformer-model scores, it also shows that this is not the case for Sentiment Projection, which performs slightly better on the original (Danish) sentences.
Finally, the workflow presented in Figure 2 has been used to design a sentiment model, but allows easy generalization to other concepts of choice. The method could also work for other emotional concepts, such as emotion recognition, language detection, or abstract concepts like a nature-to-industry gradient. We encourage curious readers to search for inspiration for potential vectors in Linear Representation Hypothesis [Peterson-2020-Parallelograms-Revisited] and Steering Vector [wehner_taxonomy_2025] literature. Due to the flexible nature of the algorithm, there is no rigid lower boundary on the number of training points required for a stable vector, although the chances of over-representing the non-concept context of training sentences naturally increase as the number of sentences decreases. A future empirical investigation of the stability of the vector when using smaller training sets would be useful.